OpenStudy (anonymous):

You earn $17.50/hr and work 40 hr/wk. Your deductions are FICA (7.65%), federal tax withholding (12.3%), and state tax withholding (6.2%). Your housing and fixed expenses are 30% of your realized income per month. You want to save 5 months' worth in an emergency fund within a year. How much do you need to save per month to fund the emergency fund, and how much discretionary money remains per month?

OpenStudy (anonymous):

dude please help me, il like pay you pleasee

OpenStudy (btaylor):

Assuming 4 weeks in a month, you make 17.50 x 40 x 4 = $2800 a month. Your deductions (7.65% + 12.3% + 6.2% + 30%) total to 56.15% of your monthly earnings. So you are left with [2800-(.5615 x 2800) = $1227.80. 5 months would be $6139. So each month, you need to save (6139 / 12 = ) $511.58. So each month, you are left with (1227.80 - 511.58 = ) $716.22 for discretionary.